# Variability in Footpoint Mapping of Bursty Bulk Flows Using Tsyganenko Models:

### v2

@@ 3.3 Relation Between Footpoint Positions and Input Parameters
In terms of MLT, the strongest correlation was observed with the BBF’s $Y_{\mathrm{GSM}}$ position (see Figure 4), with Pearson correlation coefficients around $-0.9$ in all cases. As expected, the linear coefficients indicate that on average pre-midnight events are associated with $Y_{\mathrm{GSM}}>0$, while post-midnight events correspond to $Y_{\mathrm{GSM}}<0$. Additionally, the T96 and T04 models that use $B_{\mathrm{y,IMF}}$, $B_{\mathrm{z,IMF}}$ as input parameters show a relation between MLT and the IMF clock angle ($\theta_{c}=\rm{atan}(B_{y}/B_{z})$). For instance, in the northern hemisphere if $\theta_{c}\approx 90^{\circ}$ ($B_{y,\mathrm{GSM}}>0$, pointing towards dusk) the MLT is earlier compared to $\theta_{c}\approx 270^{\circ}$ ($B_{y,\mathrm{GSM}}<0$, pointing towards dawn). An opposite behavior is observed in southern hemisphere. However, for the T01 model, this correlation is absent, which may be attributed to the differences in how the models handle IMF components.
@@
The relationship between MLAT and the input parameters of the models is more complex. As expected, cross-correlation analysis between the input variables and MLAT reveals a moderate correlation with the distance of the BBF from Earth, indicating that the MLAT shifts closer to the pole as the distance from Earth increases. Other parameters that influence MLAT vary depending on the specific model used. The Pearson and Spearman correlation coefficients, $r_{P}$ and $r_{S}$, for each model and hemisphere are presented in Figure 5. The correlation coefficient values displayed in blue indicate the correlation between the geomagnetic-related variable and distance to Earth, while the black color indicates the correlation between MLAT and distance to Earth.
For the simplest model (T89), BBFs detected at similar distances from Earth exhibit footpoints that reach lower magnetic latitudes as Hp30 values increase. A similar behavior is observed with T96 but with the coupling function $E_{KLV}=V^{4/3}B_{T}\mathrm{sin}^{2}(\theta_{c}/2)p^{1/6}$ proposed by Vasyliunas et al. (1982). This coupling function relates all the solar wind input parameters and presents larger correlation than the Sym-h index ($r_{p}=-0.18$ and $r_{s}=-0.08$). In the case of T01, the correlation with distance is more evident but no relation with input parameters was found maybe due to the low number of events. Finally, T04 presents six input parameters that estimate the exchange of energy from the interplanetary medium to the magnetosphere, the parameter with the largest correlation was $W_{1}$ (see Qin et al. (2007) for the derivation of the parameter).
Figure 5: Relation between the BBF distance to Earth and MLAT for each Tsyganenko model. The color scale indicates the input variable or derived variable from the input parameters, with the largest correlation with MLAT. Panels (a-d) correspond to the Northern Hemisphere, and panels (e-h) correspond to the Southern Hemisphere. Each panel also displays the Pearson and Spearman correlation coefficients for MLAT vs. distance represented by the x-axis (indicated in black) and MLAT vs. the variable represented by the color bar (indicated in blue).
@@ 3.4 Analysis of Model Differences Based on Input Parameters
The MLT difference between T89 and (T96, T01 and T04) as a function of $Y_{\mathrm{GSM}}$ and clock angle $\theta_{c}$ are shown in Figure 6. In the northern hemisphere, the T89 model tends to display footpoints closer to midnight compared to T96, while the opposite is observed in the southern hemisphere. This behavior is not observed when comparing to the other models (T01 and T04). The analysis of the clock angle dispersion around the mean error shows that when $\theta_{c}=90^{\circ}$ (blue dots) the MLT is later for T89 than T96, and earlier for $\theta_{c}=270^{\circ}$ (red dots). An opposite behavior is observed in the southern hemisphere. Despite the low statistics, similar trends are observed in the comparison between T89 and (T01, T04).
Figure 6: Difference between T89 and (T96, T01 and T04) against the $Y_{\mathrm{GSM}}$ position of the BBF. The color scale is interplanetary magnetic field clock angle ($\theta_{c}$) and the solid black line is the mean value.
Figure 7 shows the MLAT difference between T89 and (T96, T01 and T04) as a function of the distance of the BBF from Earth and the relative magnitudes of the proxies for Earth’s magnetospheric dynamics. This ratio indicates how the input parameter for T89, as represented by Hp30, relates to $E_{KLV}$, Sym-h, and $W_{1}$ for models T96, T01 and T04, respectively, after normalizing by their median values. In the northern hemisphere, for BBFs located at the same $r_{\mathrm{GSM}}$, when Hp30 is greater than $E_{KLV}$ (red dots), T89 footpoints are at lower latitudes compared to T96 (Figure 7a). Conversely, when $E_{KLV}$ is greater than Hp30 (blue dots) T89 footpoints are closer to the pole than the ones computed with T96. The same behavior is observed in the southern hemisphere Figure 7d. A similar trend is observed when comparing T89 with T04, although it is less apparent with T01, possibly due to lower statistics.
Figure 7: Difference in MLAT between T89 and (T96, T01 and T04) against the $r_{\mathrm{GSM}}$ position of the BBF. The color scale represents the ratio of the normalized Hp30 geomagnetic index (input parameter in T89) to the normalized input parameters with the highest correlation ($E_{KLV}$, Sym-h, and $W_{1}$ for T96, T01, and T04 respectively). The solid black line indicates the mean value. Normalization was performed using the median value.
@@ 3.5 Footpoints Proximity as a Measure of Mapping Quality
The difference between the footpoints computed using the T89 and T96 models (Figure 6a,d and Figure 7a,d) shows that more than 90% of the BBF footpoints in both hemispheres have an absolute MLT error less than $1~{}\mathrm{hour}$ and an absolute MLAT error less than $2^{\circ}$. At latitude of $70^{\circ}$, these correspond to a east-west extent of approximately $568~{}\mathrm{km}$ and a north-south extent of $222~{}\mathrm{km}$, respectively. These differences are significantly larger than the estimated mapped mean dimensions of dipolarization flux bundles (DFBs), which are typically associated with BBFs. According to Engebretson et al. (2024), a DFB with mean radius of $0.9R_{E}$ (Liu et al., 2013) mapped to the ionosphere, has an east-west extent of approximately $180~{}\mathrm{km}$ ($0.3~{}\mathrm{MLT}$ at $70^{\circ}\mathrm{N}$) and a north-south extent of about $90~{}\mathrm{km}$ ($0.81^{\circ}$). This is consistent with our estimation of BBF footprint size, detailed in Appendix B. Assuming a dipole field model and a circular cross section, we calculated a BBF cross-section in the ionosphere of approximately $110~{}\mathrm{km}\pm 30\mathrm{km}$.
Due to both statistical and physical considerations of the footpoint distances, we used two different criteria to assign a quality metric to the mapping. Based on the statistical difference between footpoints calculated with T89 and T96 models, we define a maximum distance between footpoints of ±$\pm 1$ hour in MLT and ±$\pm 2^{\circ}$ in MLAT. Meanwhile, considering the physical size of the footprint, the maximum distance is 0.3 MLT and $0.81^{\circ}$. We introduce two footpoint quality metrics for each BBF mapped event: $Q_{fp}^{\rm{stat}}$, based on the statistical criterion, and $Q_{fp}^{\rm{phys}}$, based on the physical footprint size. In both cases, $Q_{fp}=1$ indicates that the footpoint distances between T89 and T96 fall within the defined thresholds for both coordinates; $Q_{fp}=2$ indicates that only one coordinate is within threshold; and $Q_{fp}=3$ indicates that both coordinates fall outside the defined limits. In the first case, for the statistical criterion, the fraction of events with $Q_{fp}^{\rm{stat}}=\{1,2,3\}$ are 94%, 5.3% and 0.7%, respectively, for the Northern hemisphere, and 89.85%, 10.14%, and 0.05%, respectively, for the Southern hemisphere. In the second case, where the BBF footprint size is considered, the results for $Q_{fp}^{\rm{phys}}=\{1,2,3\}$ are $63.5\%$, $35.2\%$ and $1.3\%$ for the Northern hemisphere, and $46.3\%$, $53.1\%$, and $0.6\%$ for the Southern hemisphere, respectively.
@@ 4 BBF and Swarm conjunctions
In the previous section, we observed that the BBF footpoint locations predicted by different Tsyganenko models exhibit variations and that these discrepancies are influenced by geomagnetic conditions and the specific position where the BBF was detected. Consequently, defining conjunctions can be challenging due to these variations in model predictions. In this section we want to implement the analysis we have discussed to investigate the impact on MMS-Swarm BBF conjunctions.
@@
To address this, we will analyze how the number of detected conjunctions changes with different definitions of conjunction criteria. Specifically, we will investigate the impact of varying the MLAT error ($Err_{\mathrm{MLAT}}$) and the time interval between the detection of BBFs and the Swarm satellites’ path ($\Delta t$). The MLT error is set at ±$\pm 1~{}\mathrm{hour}$ because it correspond to the typical dispersion of BBF footpoints for different solar wind and magnetospheric conditions. This analysis aims to quantify how adjustments in these parameters influence the identification of conjunctions.
@@
We define a conjunction as occurring when the Swarm trace intersects the area defined by the BBF footpoint mapped to the ionosphere using the T89 and/or T96 models, including their uncertainty margins. To evaluate the quality of these conjunctions, we introduce a conjunction quality metric ($Q_{cj}$). If the Swarm trace intersects both the T89- and T96- derived footpoints, the conjunction is assigned the highest quality ($Q_{cj}=1$); if it intersects only one of the two footpoints (either from T89 or T96), the conjunction quality is assigned $Q_{cj}=2$. The footpoint in AACGM is calculated at altitude $462~{}\mathrm{km}$ for Swarm A and C, and $510~{}\mathrm{km}$ for Swarm B. Figure 8 shows an example of Swarm A, B and C traces during $\Delta t=15~{}\mathrm{minutes}$, across the south and north hemispheres, the corresponding FAC magnitude, and the BBF footpoint with the different Tsyganenko models. The shaded area of the footpoint is given by $Err_{\mathrm{MLAT}}=\pm 1^{\circ}$, and $Err_{\mathrm{MLT}}=\pm 1$ hour. In this case we detect a conjunction in the northern hemisphere with Swarm A/C.
@@
Figure 8: Example of BBF footpoint-Swarm conjunction for Swarm trace $\Delta t=15$ minutes after end time of BBF detection. (Left) Swarm A, B and C FAC measurements, and the BBF duration indicated with the two vertical dashed lines. (Right) Polar plot with Swarm traces across the North and South hemispheres and the corresponding FAC intensity. The BBF footpoint for the different Tsyganenko models are represented by circles, with shaded areas indicating the errors $Err_{MLAT}=\pm 1^{\circ}$, $Err_{MLT}=\pm 1$. The $Q_{fp}$ and $Q_{cj}$ indicate the footpoint quality and conjunction quality with Swarm A, B, and C.
@@
Figure 9 displays the number of Swarm-BBF footpoint conjunctions in the northern hemisphere and the fraction of $Q_{1}$ (i.e., $Q_{fp}=1$ and $Q_{cj}=1$, indicating that the footpoints computed by the T89 and T96 models are in close alignment, with the Swarm satellite passing through both footpoints) for two different criteria. The range for $\Delta t$ is estimated from the BBF database by assuming that the FACs are carried by Alfvén waves (Keiling, 2009). The median length of the magnetic field line from the BBF position to the northward ionospheric footpoint is 25 $R_{E}$, and the median Alfvén velocity of BBFs is $470~{}\mathrm{km~{}s^{-1}}$. Assuming that the velocity is constant, the median time delay is $\Delta t=6~{}\mathrm{min}$. When considering the 10th and 90th percentiles of the field line length and Alfvén velocity the extreme $\Delta t$ values are 1.9 and 24.5 minutes. The latitude range was selected up to ±$\pm 5^{\circ}$, which corresponds to a distance comparable to 1 hour MLT.
@@
The results reveal that Swarm B shows a greater number of conjunctions compared to Swarm A and C. As expected, the number of conjunctions increases when increasing MLAT and time interval. In general, extending the time window by 1 minute leads to a more conjunctions than adjusting the MLAT error by ±$\pm 0.5^{\circ}$. From the two lower panels, we see that the fraction of $Q=1$ conjunctions with Swarm B is more homogeneous than Swarm A and C, which present a clear region of low fraction of $Q=1$ around $\Delta t=[0,6]$ minutes and $Err_{MLAT}$ around $3^{\circ}\mathrm{N}$. When considering the BBF footprint size, we notice that in all cases less than half of the total number of conjunctions present $Q=1$. This highlight the challenges in achieving precise spatial alignment between BBF footpoints and satellite traces. Depending on the aim of the research, different approaches may be appropriate for selecting conjunctions. If one is interested in detailed case studies, it is possible to work with very close conjunctions. However, if the goal is to conduct a statistical study, increasing the allowed MLAT error and extending the time window may be necessary to capture a broader range of conjunctions.
@@
Figure 9: First row: Number of BBF footpoint-Swarm conjunctions for different time and MLAT position error ($\Delta t$, ±$\pm Err_{\mathrm{MLAT}}$ respectively) for the Northern hemisphere and each Swarm satellite. The number inside each bin is the total number of conjunctions and the color is the percentage with respect to the total number of BBFs. Second row: Percentage of conjunctions with $Q_{fp}^{\rm{stat}}=1$ (the footpoint distance computed with T89 and T96 models is lower ±$\pm 1$ hour in MLT and ±$\pm 2^{\circ}$ in MLAT) and $Q_{cj}=1$ (Swarm satellite passing through both footpoints). Third row: Similar to second row but with $Q_{fp}^{\rm{phys}}=1$ (the footpoint distance computed with T89 and T96 models is lower ±$\pm 0.3$ hour in MLT and ±$\pm 0.81^{\circ}$ in MLAT). Column (a) BBF conjunctions with Swarm A. Column (b) BBF conjunctions with Swarm B. Column (c) BBF conjunctions with Swarm C.

### v3

@@ 3.3 Relation between footpoint positions and input parameters
In terms of MLT, the strongest correlation was observed with the BBF position $Y_{\mathrm{GSM}}$ (see Figure 4), with Pearson and Spearman correlation coefficients around $-0.9$ in all cases. As expected, the linear coefficients indicate that on average pre-midnight events are associated with $Y_{\mathrm{GSM}}>0$, while post-midnight events correspond to $Y_{\mathrm{GSM}}<0$. Furthermore, the T96, T04, TA15b, and TA15n models that use $B_{\mathrm{y,IMF}}$, $B_{\mathrm{z,IMF}}$ as input parameters show a relation between MLT and the IMF clock angle ($\theta_{c}=\rm{atan}(B_{y}/B_{z})$). For example, in the Northern Hemisphere, if $\theta_{c}\approx 90^{\circ}$ ($B_{y,\mathrm{GSM}}>0$, pointing toward dusk), the MLT is earlier compared to $\theta_{c}\approx 270^{\circ}$ ($B_{y,\mathrm{GSM}}<0$, pointing toward dawn). The opposite behavior is observed in the Southern Hemisphere. However, for the T01 model, this correlation is absent, which may be attributed to differences in the way models handle IMF components.
@@
The relationship between MLAT and the input parameters of the models is more complex. As expected, cross-correlation analysis between the input variables and MLAT reveals a moderate correlation with the distance of the BBF from Earth, indicating that the MLAT shifts closer to the pole as the distance from Earth increases. Other parameters that influence MLAT vary depending on the specific model used. The Pearson and Spearman correlation coefficients between MLAT and the distance to Earth, $r_{P}$ and $r_{S}$, for each model and hemisphere are presented in Figure 5.
For the simplest model (T89), BBFs detected at similar distances from Earth exhibit footpoints that reach lower magnetic latitudes as the Hp30 values increase. A similar behavior is observed with T96 but with the coupling function $E_{KLV}=V^{4/3}B_{T}\mathrm{sin}^{2}(\theta_{c}/2)p^{1/6}$ proposed by Vasyliunas et al. (1982). This coupling function relates all the solar wind input parameters and presents a larger correlation than the Sym-h index ($r_{p}=-0.18$ and $r_{s}=-0.08$). In the case of T01, the correlation with distance is also evident, but no clear relationship with input parameters was found, possibly due to the limited number of events. In contrast, the T04, TA15b, and TA15n models exhibit weaker correlation with Earth’s distance. Furthermore, the TA15 models show no significant relationship with either the B-index or the N-index.
Figure 5: Relation between the BBF distance to Earth and MLAT for each Tsyganenko model. The color scale indicates the input variable or derived variable from the input parameters, with the largest correlation with MLAT. Panels (a-d) correspond to the Northern Hemisphere, and panels (e-h) correspond to the Southern Hemisphere. Each panel also displays the Pearson and Spearman correlation coefficients for MLAT vs. distance represented by the x-axis. The solid black line is the linear regression, while the dashed red and pink lines are the linear regression shifted ±$\pm 1^{\circ}$ and ±$\pm 5^{\circ}$ respectively.
@@ 3.4 Analysis of model differences based on input parameters
The MLT difference between TA15n and (T89, T96, T01, T04, and TA15b) as a function of $Y_{GSM}$ and the clock angle $\theta_{c}$ is shown in Figure 6. In the Northern Hemisphere, the T89 model tends to display footpoints closer to midnight compared to TA15n, while the opposite is observed in the Southern Hemisphere. This behavior is not clearly observed when comparing to the other models (T96, T01, T04, TA15b). The analysis of the dispersion of the clock angle around the mean error shows that when $\theta_{c}=90^{\circ}$ (blue dots) the MLT mapped by T89 is later than that of TA15n, and earlier for $\theta_{c}=270^{\circ}$ (red dots). An opposite trend is observed when comparing TA15n with TA15b, although the differences are smaller. A similar pattern appears in the comparison between TA15n and T04, though limited by low statistical coverage. In the Southern Hemisphere, the behavior is reversed to that observed in the Northern Hemisphere.
Figure 6: MLT difference between TA15n and (T89, T96, T01, T04, and TA15b) against the $Y_{\mathrm{GSM}}$ position of the BBF. The color scale is interplanetary magnetic field clock angle ($\theta_{c}$) and the solid black line is the mean value.
While previous results focused on comparisons with TA15n, the following analysis uses T89 as the reference model, since TA15n shows no clear relationship with Earth distance or magnetospheric activity proxies, unlike the older models, as shown in Figure 5. Accordingly, Figure 7 presents the MLAT differences between T89 and (T96, T01 and T04, TA15b, and TA15n) as a function of the distance of the BBF from Earth and the relative magnitudes of the proxies for Earth’s magnetospheric dynamics. This ratio indicates how the input parameter for T89, as represented by Hp30, relates to $E_{KLV}$, Sym-h, $W_{1}$, B-index, and N-Index for the models T96, T01, T04, TA15b, and TA15n respectively, after normalizing by their median values. In the Northern Hemisphere, for BBFs located at the same $r_{\mathrm{GSM}}$, when Hp30 is greater than $E_{KLV}$ (red dots), T89 footpoints are at lower latitudes compared to T96 (Figure 7a). Conversely, when $E_{KLV}$ is greater than Hp30 (blue dots), T89 footpoints are closer to the pole than the ones computed with T96. The same behavior is observed in the Southern Hemisphere Figure 7f. A similar trend is observed when comparing T89 with T04, although it is less apparent with T01, possibly due to lower statistics. As mentioned earlier, no clear trend is observed in the comparisons involving the TA15 models.
Figure 7: MLAT difference between T89 and (T96, T01, T04, TA15b, and TA15n) against the $r_{\mathrm{GSM}}$ position of the BBF. The color scale represents the ratio of the normalized Hp30 geomagnetic index (input parameter in T89) to the normalized input parameters with the highest correlation ($E_{KLV}$, Sym-h, $W_{1}$, B-index, and N-index for T96, T01, T04, TA15b, and TA15n respectively). The solid black line indicates the mean value. Normalization was performed using the median value.
@@ 3.5 Footpoints proximity as a measure of mapping quality
The difference between the footpoints, computed with the T89 and TA15n models (Figure 6a,f and Figure 7a,f) shows that more than 80% of the BBF footpoints in both hemispheres have an absolute MLT error less than $1~{}\mathrm{hour}$ and an absolute MLAT error less than $4^{\circ}$. At $70^{\circ}$ latitude, these correspond to an east-west extent of approximately $568~{}\mathrm{km}$ and a north-south extent of $444~{}\mathrm{km}$, respectively. These differences are significantly larger than the estimated mapped mean dimensions of dipolarization flux bundles (DFBs), which are typically associated with BBFs. According to Engebretson et al. (2024), a DFB with a mean radius of $0.9R_{E}$ (Liu et al., 2013) mapped to the ionosphere has an east-west extent of approximately $180~{}\mathrm{km}$ ($0.3~{}\mathrm{MLT}$ at $70^{\circ}\mathrm{N}$) and a north-south extent of about $90~{}\mathrm{km}$ ($0.81^{\circ}$). This is consistent with our estimation of the BBF footprint size, as detailed in the Appendix B. Assuming a dipole field model and a circular cross section, we calculated a BBF cross-section in the ionosphere of approximately $110~{}\mathrm{km}\pm 30\mathrm{km}$.
Due to both statistical and physical considerations of the footpoint distances, we used two different criteria to assign a quality metric to the mapping. Based on the statistical difference between the footpoints calculated with the T89 and TA15n models, we define a maximum distance between footpoints of ±$\pm 1$ hour in MLT and ±$\pm 4^{\circ}$ in MLAT. Meanwhile, considering the physical size of the footprint, the maximum distance is 0.3 MLT and $0.81^{\circ}$. We introduce two footpoint quality metrics for each BBF mapped event: $Q_{fp}^{\rm{stat}}$, based on the statistical criterion, and $Q_{fp}^{\rm{phys}}$, based on the physical footprint size. In both cases, $Q_{fp}=1$ indicates that the footpoint distances between T89 and TA15n fall within the defined thresholds for both coordinates; $Q_{fp}=2$ indicates that only one coordinate is within threshold; and $Q_{fp}=3$ indicates that both coordinates fall outside the defined limits. In the first case, for the statistical criterion, the fraction of events with $Q_{fp}^{\rm{stat}}=\{1,2,3\}$ are 89.0%, 9.6% and 1.4%, respectively, for the Northern Hemisphere, and 79.5%, 18.7%, and 1.8%, respectively, for the southern hemisphere. In the second case, where the BBF footprint size is considered, the results for $Q_{fp}^{\rm{phys}}=\{1,2,3\}$ are $23.8\%$, $47.7\%$ and $28.5\%$ for the Northern Hemisphere, and $14.6\%$, $51.8\%$, and $33.6\%$ for the Southern Hemisphere, respectively.
@@ 4 BBF and Swarm conjunctions
In the previous section, we observed that the BBF footpoint locations predicted by different Tsyganenko models exhibit variations and that these discrepancies are influenced by geomagnetic conditions and the specific position where the BBF was detected. Consequently, defining conjunctions can be challenging due to these variations in model predictions. In this section, we want to implement the analysis we have discussed to investigate the impact on MMS-Swarm BBF conjunctions.
@@
To address this, we will analyze how the number of detected conjunctions changes with different definitions of conjunction criteria. Specifically, we will investigate the impact of varying the MLAT error ($Err_{\mathrm{MLAT}}$) and the time interval between the detection of BBFs and the Swarm satellites’ paths ($\Delta t$). The MLT error is set at ±$\pm 1~{}\mathrm{hour}$ because it corresponds to the typical dispersion of the BBF footpoints for different solar wind and magnetospheric conditions. This analysis aims to quantify how adjustments in these parameters influence the identification of conjunctions.
@@
We define a conjunction as occurring when the Swarm trace intersects the area defined by the BBF footpoint mapped to the ionosphere using any model, including their uncertainty margins. To evaluate the quality of these conjunctions, we introduce a conjunction quality metric ($Q_{cj}$). If the Swarm trace intersects two different Tsyganenko derived footpoints, the conjunction is assigned the highest quality ($Q_{cj}=1$); if it intersects only one of the footpoints, the conjunction quality is assigned $Q_{cj}=2$. Figure 8 shows an example of Swarm A, B and C traces during $\Delta t=15~{}\mathrm{minutes}$, across the Southern and Northern Hemispheres, the corresponding FAC magnitude, and the BBF footpoint with the different Tsyganenko models. The shaded area of the footpoint is given by $Err_{\mathrm{MLAT}}=\pm 1^{\circ}$, and $Err_{\mathrm{MLT}}=\pm 1$ hour. In this example, we detect a conjunction in the Northern Hemisphere with Swarm A/C.
@@
Figure 8: Example of BBF footpoint-Swarm conjunction for Swarm trace $\Delta t=15$ minutes after end time of BBF detection. (Left) Swarm A, B and C FAC measurements, and the BBF duration indicated with the two vertical dashed lines. (Right) Polar plot with Swarm traces across the Northern and Southern Hemispheres and the corresponding FAC intensity. The BBF footpoint for the different Tsyganenko models are represented by circles, with shaded areas indicating the errors $Err_{MLAT}=\pm 1^{\circ}$, $Err_{MLT}=\pm 1$ h. The $Q_{fp}$ and $Q_{cj}$ indicate the footpoint quality and conjunction quality with Swarm A, B, and C.
@@
Figure 9 displays the number of Swarm-BBF footpoint conjunctions in the Northern Hemisphere and the fraction of $Q_{1}$ (i.e., $Q_{fp}=1$ and $Q_{cj}=1$, indicating that the footpoints computed by the T89 and TA15n models are in close alignment, with the Swarm satellite passing through two different footpoints) for two different criteria. The range for $\Delta t$ is estimated from the BBF database by assuming that the FACs are carried by Alfvén waves (Keiling, 2009). The median length of the magnetic field line from the BBF position to the northward ionospheric footpoint is 25 $R_{E}$, and the median Alfvén velocity of the BBFs is $470~{}\mathrm{km~{}s^{-1}}$. Assuming that the velocity is constant, the median time delay is $\Delta t=6~{}\mathrm{min}$. When considering the 10th and 90th percentiles of the field line length and Alfvén velocity the extreme $\Delta t$ values are 1.9 and 24.5 minutes. The latitude range was selected up to ±$\pm 5^{\circ}$, which corresponds to a distance comparable to 1 hour MLT.
@@
The results reveal that Swarm B shows a greater number of conjunctions compared to Swarm A and C. As expected, the number of conjunctions increases as the MLAT and time interval increases. In general, extending the time window by 1 minute results in a higher number of conjunctions than adjusting the MLAT error by ±$\pm 0.5^{\circ}$. From the two lower panels, we see that the fraction of $Q=1$ conjunctions with Swarm B is more homogeneous than Swarm A and C, which present a clear region of low fraction of $Q=1$ around $\Delta t=[0,6]$ minutes. When considering the size of the BBF footprint, we notice that in all cases less than half of the total number of conjunctions present $Q=1$. This highlights the challenges in achieving precise spatial alignment between BBF footpoints and satellite traces. Depending on the purpose of the research, different approaches may be appropriate for selecting conjunctions. If one is interested in detailed case studies, it is possible to work with very close conjunctions. However, if the goal is to conduct a statistical study, increasing the allowed MLAT error and extending the time window may be necessary to capture a broader range of conjunctions.
@@
Figure 9: First row: Number of BBF footpoint-Swarm conjunctions for different time and MLAT position error ($\Delta t$, ±$\pm Err_{\mathrm{MLAT}}$ respectively) for the Northern Hemisphere and each Swarm satellite. The number inside each bin is the total number of conjunctions and the color is the percentage with respect to the total number of BBFs. Second row: Percentage of conjunctions with $Q_{fp}^{\rm{stat}}=1$ (the footpoint distance computed with the T89 and TA15n models is lower ±$\pm 1$ hour in MLT and ±$\pm 4^{\circ}$ in MLAT) and $Q_{cj}=1$ (Swarm satellite passing through two different footpoints). Third row: Similar to second row but with $Q_{fp}^{\rm{phys}}=1$ (the footpoint distance computed with the T89 and TA15n models is lower ±$\pm 0.3$ hour in MLT and ±$\pm 0.81^{\circ}$ in MLAT). Column (a) BBF conjunctions with Swarm A. Column (b) BBF conjunctions with Swarm B. Column (c) BBF conjunctions with Swarm C.